# Deconstructing the Origins of Interfacial Catalysis: Why Electric Fields are Inseparable from Solvation

### v1

@@ I Introduction
The structural, dynamical and dielectric properties of aqueous interfaces have been implicated for a wide range of problems ranging from atmospheric chemistrylimmer2024 ; jungwirth2006 to the origins of lifevaida2021prebiotic ; lopezfranciscoreview2020 . Over the last decade or so there have been a flurry of studies suggesting that chemical reactions can be significantly accelerated near water interfaces compared to when they occur in the bulklopezfranciscoreview2020 . Sharpless and co-workers coined the term ‘on-water’ catalysis observing that some organic chemistry reactions dramatically speeded up when conducted in water suspensions compared to organic solvents as well as in neat conditionssharpless2005 . The consistent experimental observation is that moving from bulk water to microdroplet conditions, leads to the speed of certain chemical reactionszare2018 ; zareangewandte2015 ; cooks2011 ; Lee_Banerjee_Nam_Zare_2015 ; gordon2025 . This enhancement is thought to occur at the boundary between water and air where the chemical reactants may experience different underlying physical driving forces. The molecular origins that lead to the enhanced catalysis remains a topic of lively debate in the literature.
@@
There have been several factors that are currently being examined in the community to rationalize the origins of the enhanced chemical reactivity including the presence of large electric fields and their fluctuations (ranging between $10^{9}-10^{10}$ V/m)field12023 ; field22020 ; martinscosta2023 ; xiong2020 ; Hao2022 , interfacial solvation and anisotropy of chemical reactants induced by the air-water boundarymartins2023electrostatics ; shibdas2023 ; cookssolvation2020 , the surface activity of protons or hydroxide ionscolussi2021 ; BANERJEE2023117024 and finally, geometric confinement effectsLee_Banerjee_Nam_Zare_2015 ; zhang2022 . A particular reaction that has caught significant attention is the one leading to the formation of hydrogen-peroxide ($H_{2}O_{2}$) in microdroplets observed by the Zare groupzare2019 . Over the past few years, the Mishra group has shown through a series of careful and systematic studies that the high yields of $H_{2}O_{2}$ originally observed by the Zare group arise not from the presence of intrinsic fields at the interface, but rather from impurities such as ozone or dissolved oxygenmishra2021 ; mishra2022 ; mishra2024 ; eatoo2025 . The importance of dissolved oxygen as a source of both $H_{2}O_{2}$ and hydroxyl radicals has also recently been pointed out by other studiesasserghine2025 . On the other hand, the Head-Gordon group using molecular simulations, have implicated large electric fields and partial solvation as playing a key role in the creation of hydroxyl radicals at the air-water interfacegordonohradical2022 .
@@
The potential role of interfacial electric fields is partly rooted on the evidence that strong external static electric fields are sometimes capable of producing catalytic effects. In fact, experimental evidence has shown that electrostatic potential gradients can selectively enhance and catalyze Diels-Alder reactions Aragones and other industrially significant transformations Huang . Complementary insights into the effects of strong electric fields have come from quantum chemical computations on isolated molecules Shaik1 ; Shaik2 ; Shaik2016 and from extensive simulations on condensed-phase systems Cassone_ChemSci17 ; Cassone_TopCatal . These strong external artificial fields on the order of $\sim 10^{9}-10^{10}$ V/m align in magnitude with the local electric fields naturally occurring in condensed matter Saggu ; Boxer ; Fried_15 and are therefore thought to introduce catalytic activity. A recent study by Xie and coworkers showed using ab initio molecular dynamics simulations coupled with free energy calculations, that the barrier for a certain Diels-Alder reaction was actually slightly higher at the air-water inteface and that the intrinsic fields and their fluctuations did not help speed up the reactionjing2024jacs .
@@
The rate of a chemical reaction is typically determined by the slowest degree of freedom that shapes the lowest free energy path going from the reactant to the productlaioparrinello2002 . For a chemical process that occurs in solution, solvation can play an important role in, e.g., stabilizing the transition-state or facilitating charge-transfer events such as proton or electron transferhassanali2011 ; bagchi1989 ; mundy2018 . It is known that the fluctuations of molecular dipoles in liquid water can produce fields larger than $10^{10}$ V/m Geissler , whereas in aqueous solutions Chalmet_JCP2001 ; Saikally_PNAS2005 ; Ruiz-Lopez_PNAS2021 ; Cassone_JPCL23 and in presence of solvated ions Sellner ; Kathmann_check local field intensities exceeding $\sim 3\times 10^{10}$ V/m have been reported. Besides, electric field strengths on the order of $10^{9}$ V/m have been suggested to persist at the air-water interface xiong2020 ; Zare_PNAS2023 ; martins2023electrostatics . If these electric fields play a central role in chemical reactions, the timescales associated with their fluctuations must also be taken into consideration. In addition, since electric fields are ultimately rooted in charge density distributions arising from the water molecules oriented at the interface, can they really be de-coupled from solvation?
@@
In this work, we tackle the question of the interplay between electric fields and solvation using phenol as a test case. Electric fields at interfaces and how they modulate surface potentials has been a very active area of theoretical investigation using both classical and first principles approacheskathamnn2008 ; kathmann2011 . Our choice of using phenol is rooted in the fact that the phenolic chemical group is abundantly found in organic matter and plays a crucial role in the creation of aerosolshendrix2024 . Furthermore, due to its complex acid-base chemistry, it can also undergo various types of chemical reactionshendrix2024 . Tahara and co-workers recently demonstrated that the photodissociation of phenol occurs several orders of magnitude faster than in the bulkKusaka2021 . The precise origins of this effect remain unknown, although it is very likely tuned by specific details of how the ground and excited-state potential energy surfaces evolve differently for phenol in the bulk and at the interface, as recently shown by Moritamorita2022 . Our goal here is to use this system as a simple model to investigate how electric fields on the hydroxyl moiety of the phenol change moving from bulk to interfacial environments and to examine the connection between solvation and electric fields. We want to accomplish this using an agnostic, unsupervised learning approach based on the Information Imbalance (II) introduced by Laio and collaborators Glielmo_PNASNexus_2022 . This method allows us to directly comment on which set of collective variables are more informative in a given system.
@@
Using classical molecular dynamics simulations of phenol in different environmental conditions, we show that the electric fields along the hydroxyl group of the phenol are no different in bulk or interfacial conditions. In fact, in our current models, the electric field on a water molecule in the bulk is actually slightly larger than that of the phenol at the interface. For both the situations where the phenol is in the bulk and at the interface, we observe large fluctuations in the electric field. However, the timescales associated with these fluctuations are rather fast, ranging between $\sim 1-10$ ps. This suggests to us at least, that if electric field fluctuations are to play a role in speeding up chemistry, the reaction must occur on a similar timescale and its reaction axis must be properly aligned with the field direction within this time-frame. In fact, it is well-known that, for externally applied static electric fields, both the magnitude and the orientation determine the fate of a chemical reaction toward either catalysis or inhibition Shaik2016 ; Zhao_2025 . By examining the solvation structure of water around the phenol using the II test, we demonstrate that the electric field on the phenol is essentially described by the proximity and orientation of the nearby solvent molecules. These findings very closely mirror those of Lopez and co-workers who, using QM/MM simulations, examined the coupling between the field on a hydroxyl ion and solvent reorganizationmartinscosta2023 .
@@ II.1 Molecular Dynamics Simulations
We carried out classical molecular dynamics simulations of three systems: bulk water with a single phenol molecule in it (bulk), a water slab with a phenol molecule at the air/water interface (1 PHX) and a water slab with 25 phenol molecules at each air/water interface of the slab, with a total of 50 phenol molecules in the whole system (25 PHX). We chose this concentration of phenol molecules which results in a surface excess of phenol of approximately 0.7 molecules per nm2 corresponding to the lower concentration limit employed in the experimental work by Tahara on the photodissociation of phenol at the air/water interfaceKusaka2021 . Snapshots of the 3 simulated systems are presented in Figure1. This system also serves as a good limiting case for probing the affect of aggregation/clusters at the air-water interface on the electric fields.
@@
The various systems were prepared as follows. The bulk system consisted of 66054 water molecules and one phenol molecule (bulk), with a simulation box of 128 Å x 12 Å x 12 Å (obtained after NPT equilibration, see below). For the slab systems, we used a previously equilibrated water box containing 3636 water molecules with dimensions of 60 Å x 60 Å x 140 Å, as reported in Ref.Donkor_JCTC_2023 . To this slab, either one (1PHX) or 25 (25PHX) phenol molecules were added to each side of the interface.
@@
All simulations were conducted using a Langevin thermostat set to 300 K with a damping constant ($\gamma$) of 2.0 ps-1 and a time step of 2 fs. Periodic boundary conditions were used in all directions for all systems, for the interface systems the z-axis (normal to the interface) was large enough to ensure no interactions between periodic boundaries in that direction. The OPC water modelOPC ; OPC3 was employed since it has been shown to accurately predict the surface tension of waterOPC_surface_tension . Parameters for phenol were generated using the LEaP protocol within the AMBER package. Long-range electrostatic interactions were treated using the Particle Mesh Ewald (PME) method with a cutoff of 1 Å. Simulations were performed using the AMBERambertools2023 and GROMACSgmx_1995 ; gmx_2001 ; gmx_2005 ; gmx_2008 ; gmx_2013 ; gmx_2015_article ; gmx_2015_conference software package. The latter was used since it allows for more convenient postprocessing tools to extract the surface tension in order to validate the models of the interfacial systems which due to the low concentration (surface access) are basically the same as for the OPC neat air-water interface.
In this work, we calculated the electric field at the middle of the O-H bond of the phenol molecule. This was determined considering only the electric field produced by all the water molecules in the box for the system consisting of only one phenol. In the case of the system of 50 phenol molecules, the contribution of all the phenol molecules was also factored in addition to the water contribution, excluding the self contribution. The electric field was calculated according to Equation 1, as done in previous studiesReischl20022009 ; jong2018 , using the atomic point charges given by the OPC force field, used to describe the water molecules.
@@
In this equation $\boldsymbol{r}_{i}$ is the distance of the charge to the midpoint of the O-H bond and N is the total number of water molecules in the system. The field obtained at the center of the O-H bond can be subsequently projected along the O-H bond direction yielding a three component vector used for the ensuing analysis that will be described in more detail later.
@@ II.2 Information Imbalance (II)
In recent years machine learning has become an integral part of both conducting and understanding molecular simulationsjackson2023 . It is rather well appreciated that most molecular systems need to be described in terms of fluctuations occurring in high dimensional free energy landscapeslaio2021 . There is thus a lot of effort being placed into developing methods that help quantify, probe, and interpret the complexity that is inherent in molecular systems. In this particular work, we explore the relationship between electric fields and the solvation of phenol. Specifically, we re-cast this problem within the language of a statistical test known as the information imbalanceGlielmo_PNASNexus_2022 (II), which seeks to quantify the relationship between different distance measures applied to the same data. Within the context of our work here, we ask whether the electric field on the phenol can be rationalised solely, or at least mostly, by solvation coordinates. We begin by first briefly reviewing the II theory.
II is a measure to compare the information content between two distance measures $d_{A}$ and $d_{B}$, defined over a given data set. For example, within our context $d_{B}$ would be the distances in the electric field vector while $d_{A}$ will be the distances with respect to solvation degrees of freedom. Given these distances, one refers to $r^{ij}_{A}$ the distance rank of point $j$ with respect to point $i$ according to $d_{A}$. Similarly, $r^{ij}_{B}$ is the distance rank of $i$ with respect to $j$ according to distance $d_{B}$. In this way, one can say that $d_{A}$ is informative with respect to $d_{B}$ if points close according to $d_{A}$ are also close according to $d_{B}$. It is possible to say that $d_{A}$ is informative with respect to $d_{B}$ when points $ij$ that are close neighbors according to $d_{A}$ remain close also according to $d_{B}$.
@@
Recently, an extension to the II has been proposedWild_NatComm_2025 , where each feature that makes up a given distance measure is scaled by a weight, and then these weights are optimized through gradient descent. This formalism allows us to combine heterogeneous features, which might have very different units of measurement, in a straightforward manner, without needing to worry about scaling the data appropriately. This is referred to as the Differentiable Information Imbalance (DII) and it allows us to determine the most informative subset of coordinates or features to define a distance metric within a given data set. The DII is defined as:
@@
In this work, we primarily applied the DII to determine the information content of a set of solvation-related features on the electric field vector along the O-H bond of the phenol molecule. To perform a feature search and to avoid the problem of combinatorial explosion that would naturally arise if one were to test every possible feature subset, we use the forward greedy search algorithm introduced in Ref.30 coupled with the DII approach that is implemented in the Python package DADApy glielmo2022dadapy .

### v2

@@ I Introduction
The structural, dynamical and dielectric properties of aqueous interfaces have been implicated for a wide range of problems ranging from atmospheric chemistryLimmer et al. (2024); Jungwirth et al. (2006) to the origins of lifeDeal et al. (2021); Ruiz-Lopez et al. (2020). Over the last decade or so there have been a flurry of studies suggesting that chemical reactions can be significantly accelerated near water interfaces compared to when they occur in the bulkRuiz-Lopez et al. (2020). Sharpless and co-workers coined the term ‘on-water’ catalysis observing that some organic chemistry reactions dramatically speeded up when conducted in water suspensions compared to organic solvents as well as in neat conditionsNarayan et al. (2005). The consistent experimental observation is that moving from bulk water to microdroplet conditions, leads to the acceleration of certain chemical reactionsLai et al. (2018); Banerjee and Zare (2015); Girod et al. (2011); Lee et al. (2015); LaCour et al. (2025); Zheng et al. (2025). This enhancement is thought to occur at the boundary between water and air where the chemical reactants may experience different underlying physical driving forces. The molecular origins that lead to the enhanced catalysis remains a topic of lively debate in the literature.
@@
There have been several factors that are currently being examined in the community to rationalize the origins of the enhanced chemical reactivity including the presence of large electric fields and their fluctuations (ranging between $10^{9}-10^{10}$ V/m)Song et al. (2023a); Chamberlayne and Zare (2020); Martins-Costa and Ruiz-López (2023a); Xiong et al. (2020); Hao et al. (2022); Chan et al. (2025), interfacial solvation and anisotropy of chemical reactants induced by the air-water boundaryMartins-Costa and Ruiz-López (2023b); Nandy et al. (2023); Wei et al. (2020), curvature of the interface de la Puente and Laage (2024), the surface activity of protons or hydroxide ionsColussi et al. (2021); Banerjee (2023); Chan et al. (2025) and finally, geometric confinement effectsLee et al. (2015); Li et al. (2022). A particular reaction that has caught significant attention is the one leading to the formation of hydrogen-peroxide ($H_{2}O_{2}$) in microdroplets observed by the Zare groupLee et al. (2019). Over the past few years, the Mishra group has shown through a series of careful and systematic studies that the high yields of $H_{2}O_{2}$ originally observed by the Zare group arise not from the presence of intrinsic fields at the interface, but rather from impurities such as ozone or dissolved oxygenMusskopf et al. (2021); Gallo Jr. et al. (2022); Eatoo and Mishra (2024, 2025). The importance of dissolved oxygen as a source of both $H_{2}O_{2}$ and hydroxyl radicals has also recently been pointed out by other studiesAsserghine et al. (2025). On the other hand, the Head-Gordon group using molecular simulations, have implicated large electric fields and partial solvation as playing a key role in the creation of hydroxyl radicals at the air-water interfaceHeindel et al. (2022).
@@
The potential role of interfacial electric fields is partly rooted on the evidence that strong external static electric fields are sometimes capable of producing catalytic effects. In fact, experimental evidence has shown that electrostatic potential gradients can selectively enhance and catalyze Diels-Alder reactions Aragonès et al. (2016) and other industrially significant transformations Huang et al. (2019). Complementary insights into the effects of strong electric fields have come from quantum chemical computations on isolated molecules Shaik et al. (2004); Meir et al. (2010); Shaik et al. (2016) and from extensive simulations on condensed-phase systems Cassone et al. (2017, 2022). These strong external artificial fields on the order of $\sim 10^{9}-10^{10}$ V/m align in magnitude with the local electric fields naturally occurring in condensed matter Saggu et al. (2012); Chattopadhyay and Boxer (1995); Fried and Boxer (2015); Samantaray and Ramanan and are therefore thought to introduce catalytic activity. A recent study by Xie and coworkers showed using ab initio molecular dynamics simulations coupled with free energy calculations, that the barrier for a certain Diels-Alder reaction was actually slightly higher at the air-water interface and that the intrinsic fields and their fluctuations did not help speed up the reactionGong et al. (2024). In the same sense, Bonn’s group has carried out surface sensitive experiments that question the strength of electric fields and the time scale of their fluctuationsShirley et al. (2025).
@@
The rate of a chemical reaction is typically determined by the slowest degree of freedom that shapes the lowest free energy path going from the reactant to the productLaio and Parrinello (2002). For a chemical process that occurs in solution, solvation can play an important role in, e.g., stabilizing the transition-state or facilitating charge-transfer events such as proton or electron transferHassanali et al. (2011); Bagchi (1989); Roy et al. (2018). It is known that the fluctuations of molecular dipoles in liquid water can produce fields larger than $10^{10}$ V/m Geissler et al. (2001), whereas in aqueous solutions Ruiz-López (2001); Smith et al. (2005); Ruiz-López et al. (2021); Cassone et al. (2023) and in presence of solvated ions Sellner et al. (2013); Sellner and Kathmann (2014) local field intensities exceeding $\sim 3\times 10^{10}$ V/m have been reported. Besides, electric field strengths on the order of $10^{9}$ V/m have been suggested to persist at the air-water interface Xiong et al. (2020); Song et al. (2023b); Martins-Costa and Ruiz-López (2023b). If these electric fields play a central role in chemical reactions, the timescales associated with their fluctuations must also be taken into consideration. In addition, since electric fields are ultimately rooted in charge density distributions arising from the water molecules oriented at the interface, can they really be de-coupled from solvation?
@@
In this work, we tackle the question of the interplay between electric fields and solvation using phenol as a test case. Electric fields at interfaces and how they modulate surface potentials has been a very active area of theoretical investigation using both classical and first principles approachesKathmann et al. (2008, 2011). Our choice of using phenol is rooted in the fact that the phenolic chemical group is abundantly found in organic matter and plays a crucial role in the creation of aerosolsRichter et al. (2024). Furthermore, due to its complex acid-base chemistry, it can also undergo various types of chemical reactionsRichter et al. (2024). Tahara and co-workers recently demonstrated that the photodissociation of phenol occurs several orders of magnitude faster than in the bulkKusaka et al. (2021). The precise origins of this effect remain unknown, although it is very likely tuned by specific details of how the ground and excited-state potential energy surfaces evolve differently for phenol in the bulk and at the interface, as recently shown by MoritaIshiyama et al. (2022). Our goal here is to use this system as a simple model to investigate how electric fields on the hydroxyl moiety of the phenol change moving from bulk to interfacial environments and to examine the connection between solvation and electric fields. We want to accomplish this using an agnostic, unsupervised learning approach based on the Information Imbalance (II) introduced by Laio and collaborators Glielmo et al. (2022a). This method allows us to directly comment on which set of collective variables are more informative in a given system.
@@
Using classical molecular dynamics simulations of phenol in different environmental conditions, we show that the electric fields along the hydroxyl group of the phenol are no different in bulk or interfacial conditions. In fact, in our current models, the electric field on a water molecule in the bulk is actually slightly larger than that of the phenol at the interface. For both the situations where the phenol is in the bulk and at the interface, we observe large fluctuations in the electric field. However, the timescales associated with these fluctuations are rather fast, ranging between $\sim 1-10$ ps. This suggests to us at least, that if electric field fluctuations are to play a role in speeding up chemistry, the reaction must occur on a similar timescale and its reaction axis must be properly aligned with the field direction within this time-frame. In fact, it is well-known that, for externally applied static electric fields, both the magnitude and the orientation determine the fate of a chemical reaction toward either catalysis or inhibition Shaik et al. (2016); Zhao et al. (2025). By examining the solvation structure of water around the phenol using the II test, we demonstrate that the electric field on the phenol is essentially described by the proximity and orientation of the nearby solvent molecules. These findings very closely mirror those of Ruiz-López and co-workers who, using QM/MM simulations, examined the coupling between the field on a hydroxyl ion and solvent reorganizationMartins-Costa and Ruiz-López (2023a).
@@ II.1 Molecular Dynamics Simulations
We carried out classical molecular dynamics simulations of three systems: bulk water with a single phenol molecule in it (bulk), a water slab with a phenol molecule at the air/water interface (1 PHX) and a water slab with 25 phenol molecules at each air/water interface of the slab, with a total of 50 phenol molecules in the whole system (25 PHX). We chose this concentration of phenol molecules which results in a surface excess of phenol of approximately 0.7 molecules per nm2 corresponding to the lower concentration limit employed in the experimental work by Tahara on the photodissociation of phenol at the air/water interfaceKusaka et al. (2021). Snapshots of the 3 simulated systems are presented in Figure1. This system also serves as a good limiting case for probing the effect of aggregation/clusters at the air-water interface on the electric fields.
@@
The various systems were prepared as follows. The bulk system consisted of 66054 water molecules and one phenol molecule (bulk), with a simulation box of 128 Å x 12 Å x 12 Å (obtained after NPT equilibration, see below). For the slab systems, we used a previously equilibrated water box containing 3636 water molecules with dimensions of 60 Å x 60 Å x 140 Å, as reported in Ref.Donkor et al. (2023). To this slab, either one (1PHX) or 25 (25PHX) phenol molecules were added to each side of the interface.
@@
All simulations were conducted using a Langevin thermostat set to 300 K with a damping constant ($\gamma$) of 2.0 ps-1 and a time step of 2 fs. Periodic boundary conditions were used in all directions for all systems, for the interface systems the z-axis (normal to the interface) was large enough to ensure no interactions between periodic boundaries in that direction. The OPC water modelIzadi et al. (2014); Izadi and Onufriev (2016) was employed since it has been shown to accurately predict the surface tension of water. Specifically, at room temperature OPC yields a surface tension of 75 mN/m in excellent agreement with the experimentally measured 70 mN/m Tempra et al. (2022). Parameters for phenol were generated using the LEaP protocol within the AMBER package. Long-range electrostatic interactions were treated using the Particle Mesh Ewald (PME) method with a cutoff of 1 Å. Simulations were performed using the AMBERCase et al. (2023) and GROMACSBerendsen et al. (1995); Lindahl et al. (2001); Van Der Spoel et al. (2005); Hess et al. (2008); Pronk et al. (2013); Abraham et al. (2015); Páll et al. (2015) software package. The latter was used since it allows for more convenient postprocessing tools to extract the surface tension in order to validate the models of the interfacial systems which due to the low concentration (surface access) are basically the same as for the OPC neat air-water interface.
In this work, we calculated the electric field at the middle of the O-H bond of the phenol molecule. This was determined considering only the electric field produced by all the water molecules in the box for the system consisting of only one phenol. In the case of the system of 50 phenol molecules, the contribution of all the phenol molecules was also factored in addition to the water contribution, excluding the self contribution. The electric field was calculated according to Equation 1, as done in previous studiesReischl et al. (2009); Jong and Hassanali (2018), using the atomic point charges given by the OPC force field, used to describe the water molecules. Since the GAFF force fields used to describe the interaction in these systems don’t include polarization and charge transfer effects we validated this methodology using the QM/MM approach that is described in the SI text.
@@
In this equation $\boldsymbol{r}_{i}$ is the distance of the charge to the midpoint of the O-H bond and N is the total number of water molecules in the system. The field obtained at the center of the O-H bond can be subsequently projected along the O-H bond direction yielding a three component vector, according to equation 2, used for the ensuing analysis that will be described in more detail later.
$proj\boldsymbol{E}=(\boldsymbol{E}\cdot\hat{\boldsymbol{r}})\hat{\boldsymbol{r}}$ (2)
@@ II.2 Information Imbalance (II)
In recent years machine learning has become an integral part of both conducting and understanding molecular simulationsJackson et al. (2023). It is rather well appreciated that most molecular systems need to be described in terms of fluctuations occurring in high dimensional free energy landscapesGlielmo et al. (2021). There is thus a lot of effort being placed into developing methods that help quantify, probe, and interpret the complexity that is inherent in molecular systems. In this particular work, we explore the relationship between electric fields and the solvation of phenol. Specifically, we re-cast this problem within the language of a statistical test known as the information imbalanceGlielmo et al. (2022a) (II), which seeks to quantify the relationship between different distance measures applied to the same data. Within the context of our work here, we ask whether the electric field on the phenol can be rationalised solely, or at least mostly, by solvation coordinates. We begin by first briefly reviewing the II theory.
[…]
Given these distances, one refers to $r^{ij}_{A}$ the distance rank of point $j$ with respect to point $i$ according to $d_{A}$. Similarly, $r^{ij}_{B}$ is the distance rank of $i$ with respect to $j$ according to distance $d_{B}$. In this way, one can say that $d_{A}$ is informative with respect to $d_{B}$ if points close according to $d_{A}$ are also close according to $d_{B}$. It is possible to say that $d_{A}$ is informative with respect to $d_{B}$ when points $ij$ that are close neighbors according to $d_{A}$ remain close also according to $d_{B}$.
@@
Recently, an extension to the II has been proposedWild et al. (2025), where each feature that makes up a given distance measure is scaled by a weight, and then these weights are optimized through gradient descent. This formalism allows us to combine heterogeneous features, which might have very different units of measurement, in a straightforward manner, without needing to worry about scaling the data appropriately. This is referred to as the Differentiable Information Imbalance (DII) and it allows us to determine the most informative subset of coordinates or features to define a distance metric within a given data set. The DII is defined as:
@@
In this work, we primarily applied the DII to determine the information content of a set of solvation-related features on the electric field vector along the O-H bond of the phenol molecule. To perform a feature search and to avoid the problem of combinatorial explosion that would naturally arise if one were to test every possible feature subset, we use the forward greedy search algorithm introduced in Ref.62 coupled with the DII approach that is implemented in the Python package DADApy Glielmo et al. (2022b).